OpenStudy (anonymous):

A company sells boxes of cow bells (b) for $20 and boxes of air horns (h) for $30. They can make a batch of cow bells that fills 8 boxes or a batch of air horns that fills 9 boxes. The company only has 53 boxes. Which system of equations helps the company plan to make $275?

OpenStudy (anonymous):

we can set the target on one side, $275 = and also, they can't make one cow bell. they must make a batch, of each of them. so on the Right Hand Side we must have something like 8*b*$20 + 9*h*$30, because if we increase "b" by one, the only possible increase is an entire batch which comes out to be 8 bells at $20 each

OpenStudy (anonymous):

"system of equations" the other restriction is that the total boxes are 53. There could be more than one possible combination yielding $275. the second equation tells us that 8*b*+ 9*h, must be 53 or less boxes (the configuration of batches can not exceed available boxes)

OpenStudy (anonymous):

b = 53 - h 8b + 9h = 53 8b = 53 - 9h 20b + 30h = 275 b = 53 - h 20b + 30h = 275 8b = 53 - 9h 20b - 30h = 275 thoese are the pssible awnsers

OpenStudy (anonymous):

Did u ever get the answer to this ?